OpenStudy (anonymous):

im so confused with the changing of signs. :( how do you do this? factorise ac-3c+6b-2ab

OpenStudy (anonymous):

distributive property? :/

OpenStudy (anonymous):

i don't get it :( why must a = (a-3), b = 2b, c = c. where did u get all those numbers....

OpenStudy (anonymous):

okay :) hey, i did it this way but i didn't get the ans :/ ac-3c+6b-2ab =c(a-3)+2b(3-a) =c(a-3)-2b(a-3) =(c-2b)(a-3) did i changed my signs wrongly or smth? :(

OpenStudy (anonymous):

Wait...let me double check mine

OpenStudy (anonymous):

You got it correct

OpenStudy (anonymous):

I rushed my answer. Sorry. I will delete it.

OpenStudy (anonymous):

okay.. hey but i did it another way and i got this! ac-3c+6b-2ab =ac-3c-2ab+6b =c(a-3)-2b(a+3) =c(a-3)+2b(a-3) =(c+2b)(a-3) which is correct? :/

OpenStudy (anonymous):

wait ....

OpenStudy (anonymous):

Your step three is wrong. The way you did it the very first time is correct.

OpenStudy (anonymous):

why? :( how do you know? sorry, i know i ask a lot...

OpenStudy (anonymous):

In your step three, when you distribute the negative you would get (-a-3) not (a-3)

OpenStudy (anonymous):

-(a+b) = -a+-b=-a-b

OpenStudy (anonymous):

okay.. hey but i did it another way and i got this! 0. ac-3c+6b-2ab 1. =ac-3c-2ab+6b 2. =c(a-3)-2b(a+3) 3. =c(a-3)+2b(a-3) <---- Incorrect distribution of negative 4. =(c+2b)(a-3)

OpenStudy (anonymous):

:o ! actually i think my step 2 is wrong. ac-3c+6b-2ab 1. =ac-3c-2ab+6b 2. =c(a-3)-2b(a+3 ) --> this should be c(a-3)-2b(a-3) . is it correct?

OpenStudy (anonymous):

ac - 3c + 6b - 2ab c( a-3) +2b( 3 - a) c(a -3 ) -2b( a-3) (a-3) (c -2b)....i think Cherry did it right!

OpenStudy (anonymous):

c(a-3)-2b(a+3) is correct because you only factor out a positive 2b. c(a-3)-2b(a-3) is incorrect because once again, you factored out a positive 2b. You didn't factor out any negatives. Here's your step 1 re-written with parentheses: c(a-3)-(2b(a+3))

OpenStudy (anonymous):

heyhey thanks guys :) esp heromiles for trying to explain the steps to me.:) i rly apprecite it.

OpenStudy (anonymous):

You're very welcome